# ROTH-TYPE THEOREMS IN ADDTIVE COMBINATORICS

### v1

@@ 3.2. Roth’s Theorem and the Relative Version in the Primes
(b)(Large Rank)There exists $l$ l.i. columns of $\mathbf{C}$ s.t. any of them excluded, the remaining n1n-1 columns can be devided into two groups among each there are $l$ l.i columns.
@@
$\sum\limits_{n\in A}\frac{1}{n}=\int_{1-\epsilon}^{\infty}\frac{\text{d}F(t)}{t}\leqslant 1+\int_{1-\epsilon}^{\infty}\frac{F(t)}{t^{2}}\text{d}t\ll\int_{1-\epsilon}^{\infty}\frac{1}{t(\log t)^{1+c}}\text{ converges.}\qed$
@@ 4.1.3. Density Increment Strategy
1. (Random-Many 3APs) $\Lambda(A)/|B|^{2}=\langle 1_{A}\ast 1_{A},1_{2A}\rangle_{\mu_{B}}\gg_{d}\alpha^{3}$;
@@
Combining this with (5) to further bound with $\alpha$ the implicit constant w.r.t. $d_{n}$, we derive a lower bound for $\Lambda(A)$. It turns out that this strategy is rather effective and contributes to the polylogarithmic bound with large exponent and even the quasi-polynomial bound, due to Bourgain[7], Kelley-Meka[27] and Bloom-Sisask[5][6].
@@ 4.3. Polynomial Method in the Finite Field Case
He actually proved that $|K|\geqslant\binom{q+n-1}{n}=|M_{n}|$, where $M_{n}=\{\text{monomials in }\mathbb{F}_{q}[x_{1},...,x_{n}],\deg\leqslant q-1\}.$ For any field $\mathbb{F}$, by viewing polynomials of $n$ variables as functions in $\mathbb{F}^{n}$, we know the following fundamental fact.
@@ 5.1. Relax the Conditions in Multidimensional Linear Roth’s theorem
where $\mathbf{c}_{i}\in G\{0\}$ while $\mathbf{x}_{i}\in\Delta$ for all $1\leqslant i\leqslant n$, and the dot multiplication is the common inner product by coordinate. The number of solutions in $A$ to the original system then equals the number of solutions in $\Delta(A)$ to (9), which is

### v2

@@ 3.2. Roth’s Theorem and the Relative Version in the Primes
(b)(Large Rank)There exists $l$ linearly independent columns of $\mathbf{C}$ s.t. any of them excluded, the remaining n1n-1 columns can be devided into two groups among each there are $l$ linearly independent columns.
@@
$\sum\limits_{n\in A}\frac{1}{n}=\int_{1-\epsilon}^{\infty}\frac{\text{d}F(t)}{t}\leqslant 1+\int_{1-\epsilon}^{\infty}\frac{F(t)}{t^{2}}\text{d}t\ll\int_{1-\epsilon}^{\infty}\frac{\text{d}t}{t(\log t)^{1+c}}\text{ converges.}\qed$
@@ 4.1.3. Density Increment Strategy
1. (Random-Many 3-APs) $\Lambda(A)/|B|^{2}=\langle 1_{A}\ast 1_{A},1_{2A}\rangle_{\mu_{B}}\gg_{d}\alpha^{3}$;
@@
Combining this with (5) to further bound with $\alpha$ the implicit constant w.r.t. $d_{n}$, we derive a lower bound for $\Lambda(A_{n})$, thus for $\Lambda(A)$ by translation invariance. It turns out that this strategy is rather effective and contributes to the polylogarithmic bound with large exponent and even the quasi-polynomial bound, due to Bourgain[7], Kelley-Meka[27] and Bloom-Sisask[5][6].
@@ 4.3. Polynomial Method in the Finite Field Case
He actually proved that $|K|\geqslant\binom{q+n-1}{n}=|M_{n}|$, where $M_{n}=\{\text{monomials in }\mathbb{F}_{q}[x_{1},...,x_{n}],\deg\leqslant q-1\}.$ For any field $\mathbb{F}$, by viewing polynomials of $n$ variables over $\mathbb{F}$ as functions on $\mathbb{F}^{n}$, we know the following fundamental fact.
@@ 5.1. Relax the Conditions in Multidimensional Linear Roth’s theorem
where $\mathbf{c}_{i}\in G\backslash\{0\}$ while $\mathbf{x}_{i}\in\Delta$ for all $1\leqslant i\leqslant n$, and the dot multiplication is the common inner product by coordinate. The number of solutions in $A$ to the original system then equals the number of solutions in $\Delta(A)$ to (9), which is